OpenStudy (anonymous):

What is the Least Common Denominator of 4/x + 2x/x-1 + x+3/2x

OpenStudy (anonymous):

\[\frac{4}{x}+\frac{2x}{x-1}+\frac{x+3}{2x}\] lcm is \[2x(x-1)\]

OpenStudy (anonymous):

rather LCD

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw